OpenStudy (anonymous):

A salad bar offers 8 choices of toppings for lettuce. In how many ways can you choose 4 or 5 toppings?

OpenStudy (anonymous):

combinations 8C4 + 8C5 = 8*7*6*5 8*7*6 ------- + ---- 4*3*2*1 3*2*1 = 70 + 56 = 126